# Spherically symmetric solutions to the Einstein-scalar field conformal constraint equations

### v1

@@ Abstract.
Recent works [18, 43] have shown that the Einstein–scalar field conformal constraint equations are highly complex and generally intractable, even in the vacuum case. In this article, to gain a clearer understanding and offer a new perspective, we study these equations under special assumptions: the manifold $(M,g)$ is harmonic and data is radial.
In this setting, the system reduces to a single nonlinear equation and is completely resolved in the standard cases. In particular, on the sphere, our results reveal phenomena that contrast with the well-known achievements on compact manifolds without conformal Killing vector fields, including nonexistence of solutions in the near-CMC regime and instability when the mean curvature is non-constant. By contrast, on Euclidean or hyperbolic manifolds, the equations are always solvable, with all expected properties of solutions satisfied. These findings support the view that, although the conformal method appears to present some drawbacks on compact manifolds, it remains a promising tool for parametrizing solutions to the constraint equations on asymptotically flat and hyperbolic manifolds in arbitrary mean curvature regimes.
In this article, we also investigate the sign of mass, showing that the ADM and asymptotically hyperbolic mass can take arbitrary sign when the decay rate of symmetric $(0,2)$-tensor $k$ at infinity is critical. Finally, most solution classes in our framework are explicit, providing a variety of models in general relativity and offering insights into the structure and behavior of initial data, particularly in numerical applications.
@@ 1. Introduction
With this framework in place, many interesting results can be established. Before presenting them, however, we first introduce the notation used in the article. From now on, $r$ is the distance function to a fixed point and we denote by $f^{\prime}$ the derivative of $f$ with respect to $r$. We remark also that, since the equations involve finding $(\varphi,W)$, and one can be recovered once the other is known, for simplicity of expression, we will sometimes say that $(\varphi,W)$ or simply $\varphi$ is a solution of (1.2). Now, let $M_{\kappa}$ be the space form of curvature $\kappa\in\{1,0,-1\}$, that is $M_{\kappa}=\mathbb{S}^{n}$ if $\kappa=1$, $M_{\kappa}=\mathbb{R}^{n}$ if $\kappa=0$ and $M_{\kappa}=\mathbb{H}^{n}$ if $\kappa=-1$. The sine and cosine function $\mathrm{s}_{\kappa}$ and $\mathrm{c}_{\kappa}$ are
@@
This result is the source of subsequent developments in the article. In particular, point (i) provides a necessary condition for the existence of solutions, while point (ii) shows that constructing radial solutions to the conformal equations (1.2) becomes quite simple. Indeed, it suffices to choose a potential function $V$ together with radial functions $(\psi(r),\rho(r),\varphi(r))$ such that $S_{V,\psi,\rho,\varphi}>0$. Once this condition is satisfied, the identity (1.6) uniquely determines the corresponding mean curvature function $\tau(r)$, thereby completing the construction of a full data set $(V,\psi(r),\rho(r),\tau(r))$ along with a solution $\varphi(r)$ to the conformal equations (1.2). As shown in Subsection A, seeking data that satisfies this positivity condition is not difficult. Thus, our theorem yields a broad class of explicit solutions to the constraint equations, which is valuable not only as theoretical examples but also as practical models in numerical relativity, where exact solutions are scarce and highly informative.
For the further results in this article, apart from the instability phenomena on the sphere described in Main Theorem 5, we are interested in the solvability of the conformal equations (1.2) in the case where $V$, $\psi$, and $\sigma$ vanish identically, that is
@@
By studying (1.7) in this context, we provide a new perspective that, contrary to unexpected results in [43, 18], the conformal equations (1.7) in radial settings are solvable in standard spaces, particularly in the hyperbolic and Euclidean cases, with all expected properties of solutions satisfied. This give evidence that the conformal method is still an effective parameterization of solutions to the constraint equations (1.1) for a broad class of data, not only when the mean curvature is constant or almost constant. Since the geometrical structure and the related problems for the conformal equations differ on the sphere, Euclidean space, and hyperbolic space, for ease of the presentation, we present the remaining discussion and results according to the type of manifolds as follows.
@@
In the first part of this work, we are concerned with the question of the existence of solutions to the conformal equations for a given radial setting on the sphere in various mean curvature regimes. The outcomes we obtain are different from those previously known on compact manifolds without CKVF. The first distinction arises from existence/nonexistence of solutions in the near-constant mean curvature regime as we will show below. For manifolds having no CKVF, previous works such as [30, 36, 14] show that the reduced system (1.7) admits a unique solution $\varphi$ whenever $\rho\not\equiv 0$ and the ratio $d\tau\over\tau$} is sufficiently small. Solutions in this case are the so-called near-CMC solutions, and their existence may be obtained through several approaches. However, in the theorem below, we will see that this conclusion no longer holds when the manifold has CKVF, with the following setting on the round sphere as a counterexample.
@@
which coincides with the conformal equations (1.2) of null TT-tensor $\sigma$ in the vacuum case. The question of solvability of these equations was first posed by D. Maxwell in [38] and it was shown by the second author in [43, 44] to be central to the understanding of the full vacuum case of (1.2). In the absence of CKVF, the second author also proved in [43] that if $\tau>0$ satisfies
@@
  • When $\mathcal{B}_{\tau,\psi}<0$, that is the defocusing case, the system is known to be stable in the CMC and near-CMC settings thanks to well-known arguments in [3, 36, 14]. However, when $\tau$ is far–from–CMC, the stability question still remains open.
@@
For $\mathcal{B}_{\tau,\psi}<0$, we focus on a simplified setting in which $(V,\psi,\sigma)$ vanish identically, that is the equations (1.7). Using the condition $S_{V,\psi,\rho,\varphi}\geq 0$, we show that, within the space of radial functions, the equations are not only stable, but that the set of solutions is also uniformly bounded independently of the choice of $\tau$. This suggests that the stability property of the conformal equations may persist more generally in this case. When $\mathcal{B}_{\tau,\psi}$ changes sign, the situation is markedly different. In contrast to the stability result obtained by O. Druet and E. Hebey [19] for the CMC case in dimensions $n\leq 5$ as mentioned above, thanks to Identity (1.6), we will construct a sequence of blowing-up solutions to (1.2), thereby showing that the stability property of the conformal equations no longer holds in general in the non-CMC setting. This gives a negative answer to the stability question in the last case in all dimensions. More precisely, our associated results are that
Main Theorem 5 (Positive $\mathcal{B}_{\tau,\psi}$ and stability).
@@ 1.2. Hyperbolic manifold
Then the conformal equations (1.7) associated with $(\rho,\tau)$ admit at least one radial solution. Moreover, the set of radial solutions is compact.
@@ 1.3. Euclidean manifold
In this article, by restricting to radial settings, we do not encounter these limitations, as we can apply Main Theorem 1.6 to solve (1.7) on Euclidean space. Our main result is that, for any radial data set, the equations are always solvable, and moreover, the set of solutions is uniformly bounded independently of $\tau$.
@@
There exists a vacuum constraint solution $(\mathbb{R}^{n},g,k)$ such that
$g-\delta_{\text{Euc}}=O\bigg({1\over r^{n-2}}\bigg),\qquad k=O\bigg({1\over r^{n/2}}\bigg)$
at infinity and the ADM mass of $(\mathbb{R}^{n},g)$ is negative.
@@
In Section 3, provided that the manifold is harmonic, building on the result in [41], we will reduce the conformal equations (1.2) of radial settings to a single nonlinear equation involving functions of the distance, which is much easier to resolve.
In Sections 4, 5 and 6, a complete description of this single equation will be obtained in some special cases where the manifold is the sphere, the hyperbolic space or the Euclidean one. Each of these sections begins with the proof of our Main Theorem 1 in the setting considered, and then derives applications by studying the properties of solutions.
@@ 2.1. Preliminaries
There is a rich literature on the existence and classification of harmonic manifolds. The most basic examples are Euclidean spaces and the Rank One Symmetric Spaces (ROSS). In 1944, A. Lichnerowicz conjectured — proving it in dimension 4 — that these are the only harmonic manifolds. The conjecture was later confirmed by Z.I. Szabó for compact, simply connected manifolds of any dimension (see [50]). However, in a significant development, E. Damek and F. Ricci showed in [15] that the conjecture fails in the non-compact case by constructing harmonic homogeneous manifolds that are not ROSS. To date, the only known harmonic manifolds are the Euclidean spaces, ROSS, and Damek–Ricci spaces. For an introduction to the basic structure of ROSS, we refer the reader to [6, Chapter 3] or [34]. General properties of harmonic manifolds can be found in [6, chapter 6], [50, sections 1 and 2] and [33, section 2].
@@
Point (a) is a standard result of Riemannian geometry. For point (b), see [33, Corollary 2.8 and Section 6]. Point (c) follows from Szabó’s theorem and the standard properties of compact ROSS in [50]. ∎
@@ 2.2. Primary functions for the conformal equations
$u_{0}^{\prime\prime}+hu_{0}^{\prime}+\tfrac{n}{n-1}\mathrm{Ric}u_{0}=0,\quad\forall r\in(0,\mathrm{diam}(M))$ (2.5)
$u_{0}(0)=1,\,u_{0}^{\prime}(0)=0.$
@@
Assume that $(M,g)$ satisfies one of two conditions (2.8). For any $f_{v}$ given as in (2.9), we define $F_{v}$ as follows.
@@ 3.2. Reduction of the conformal equations
$V\in C^{0}(\mathbb{R}),\quad\psi\in C^{1}(M),\quad\tau\in\ C^{1}(M)\quad\text{and}\quad\rho\in C^{0}(M).$
@@ 4.1. Lichnerowicz-type equation on the sphere
By Definition (2.12) we have from (4.1)
@@ 4.2. Non-existence of solutions
In the following result, we will give a different perspective on the conformal equations on the sphere. In fact, the presence of CKVF makes the above results appear no longer valid, otherwise, it leads to an unexpected situation in which the set of solutions to the vacuum conformal equations of null $\sigma$ is infinite. To see this, we first remark that the set of all radial mean curvatures $\tau$ on the sphere well covers the non-generic inequality considered in [43, 44], that is, there exists a class of radial mean curvatures satisfying (4.13). A simple example of this is $\tau_{0}=\cos(r)+2$. By straightforward calculations, since
@@ 4.3. Existence of solutions
    $\sup_{a\in[-c,c]}\bigg(\frac{\max|(\tau_{0}-a\tau_{1})^{\prime}|}{\min|(\tau_{0}-a\tau_{1})|}\bigg)<\frac{2}{\pi}.$
@@
$\sup_{a\in[-c,c]}\Big(\frac{\max|(\tau_{0}-a\tau_{1})^{\prime}|}{\min|(\tau_{0}-a\tau_{1})|}\Big)<{2\over\pi},$
@@
has no nontrivial $W$ for all $\lambda\in(0,1]$, then the vacuum conformal equations (1.2) admit at least one solution $(\varphi,W)$. Moreover, in this case, the set of solutions to (1.2) is compact. The results in [43, 44, 18] also show that the limit equation (4.36) appears to play a central role in the solvability of the conformal method, at least for $\tau$ not changing sign. Technically, as we may see in the proof of the main result in [14] or [42], a nontrivial solution $W$ to the limit equation (4.36) arises from a blow-up phenomenon of a sequence of solutions $\{\varphi_{i}\}$, in which, after an appropriate rescaling, one obtains
$\bigg({\varphi_{i}\over\|\varphi_{i}\|_{L^{\infty}}}\bigg)^{N}\longrightarrow\sqrt{n\over n-1}{|LW|\over\tau}.$ (4.37)
Therefore, in the context of our consideration, when we restrict all to the spaces of radial functions on the sphere, that is $\varphi_{i}$ and $\tau$ in (4.37) are radial, the convergence tells us that $|LW|$ is also radial, and hence by Proposition (3.3), $W$ must be of the form $W=u(r)dr$ for some radial function $u(r)$. However, in this case, the following proposition shows that such a non-trivial solution $W$ does not exist.
@@
Although the limit equation criterion cannot be applied for the conformal equations on the sphere due to the presence of CKVF, this proposition provides insight that the equations (4.12) seems to be always solvable in some sense. In the next two results, we will give existence of solutions to (4.12) in the far–from–CMC regime by generalizing the results of Theorem 4.7, in which the conditions requiring mean curvature to be near constant and time-derivative $\rho$ to be small may be dropped; instead, all we need to assume is that $(\tau_{0}+\alpha\tau_{1})$ does not change sign or that $\rho$ is $C^{1}$-regular. More precisely, our results are as follows.
@@
an analysis similar to that in [14, Lemma 2.6] or [42, Theorem 3.3] shows that (up to a subsequence)

### v2

@@ Abstract.
Recent works [15, 36] have shown that the Einstein–scalar field conformal constraint equations are highly complex and generally intractable, even in the vacuum case. In this article, to gain a clearer understanding and offer a new perspective, we study these equations under special assumptions: the manifold $(M,g)$ is harmonic and all data are radial.
In this setting, the system reduces to a single nonlinear equation and is completely solved in the standard cases. In particular, on the sphere, our results reveal phenomena that contrast with the well-known achievements on compact manifolds without conformal Killing vector fields, including nonexistence of solutions in the near-CMC regime and instability when the mean curvature is non-constant. By contrast, on Euclidean or hyperbolic manifolds, the equations are always solvable, with all expected properties of solutions satisfied. These findings support the view that, although the conformal method appears to present some drawbacks on compact manifolds, it remains a promising tool for parametrizing solutions to the constraint equations on asymptotically flat and hyperbolic manifolds in arbitrary mean curvature regimes.
In this article, we also investigate the sign of mass, showing that the ADM and asymptotically hyperbolic mass of vacuume constraint solutions can take arbitrary sign when the decay rate of symmetric $(0,2)$-tensor $k$ at infinity is critical. Finally, most solution classes in our framework are explicit, providing a variety of models in general relativity and offering insights into the behavior of initial data, particularly in numerical applications.
@@ 1. Introduction
With this framework in place, many interesting results can be established. Before presenting them, however, we first introduce the notation used in the article. From now on, $r$ is the distance function to a fixed point and we denote by $f^{\prime}$ the derivative of $f$ with respect to $r$. We remark the the unknown of the conformal equations is a couple $(\varphi,W)$ of a function and a 1-form. However, in some situations, the 1-form $W$ can be recovered once the function $\varphi$ is known. For simplicity of expression, we will sometimes we will sometime just say that $\varphi$ is a solution of (1.2). Now, let $M_{\kappa}$ be the space form of curvature $\kappa\in\{1,0,-1\}$, that is $M_{\kappa}=\mathbb{S}^{n}$ if $\kappa=1$, $M_{\kappa}=\mathbb{R}^{n}$ if $\kappa=0$ and $M_{\kappa}=\mathbb{H}^{n}$ if $\kappa=-1$. The sine and cosine function $\mathrm{s}_{\kappa}$ and $\mathrm{c}_{\kappa}$ are
@@
This result is the source of subsequent developments in the article. In particular, point (i) provides a necessary condition for the existence of solutions, while point (ii) shows that constructing radial solutions to the conformal equations (1.2) becomes quite simple. Indeed, it suffices to choose a potential function $V$ together with radial functions $(\psi(r),\rho(r),\varphi(r))$ such that $S_{V,\psi,\rho,\varphi}>0$. Once this condition is satisfied, the identity (1.6) uniquely determines the corresponding mean curvature function $\tau(r)$, thereby completing the construction of a full data set $(V,\psi(r),\rho(r),\tau(r))$ along with a solution $\varphi(r)$ to the conformal equations (1.2). As shown in Appendix A, seeking data that satisfies this positivity condition is not difficult. Thus, our theorem yields a broad class of explicit solutions to the constraint equations, which is valuable not only as theoretical examples but also as practical models in numerical relativity, where exact solutions are scarce and highly informative.
For the further results in this article, apart from the instability phenomena on the sphere described in Main Theorem 5, we are interested in the existence of solutions to the conformal equations (1.2) in the case where $V$, $\psi$, and $\sigma$ vanish identically, that is
@@
By studying (1.7) in this context, we provide a new perspective that, contrary to unexpected results in [36, 15], the conformal equations (1.7) in radial settings are solvable in certain standard spaces, particularly in the hyperbolic and Euclidean cases, with all expected properties of solutions satisfied. This give evidence that the conformal method is still an effective parameterization of solutions to the constraint equations (1.1) for a broad class of data, not only when the mean curvature is constant or almost constant. Since the geometrical structure and the related problems for the conformal equations differ on the sphere, Euclidean space, and hyperbolic space, for ease of the presentation, we present the remaining discussion and results according to the type of manifolds as follows.
@@
In the first part of this work, we are concerned with the question of the existence of solutions to the conformal equations for a given radial setting on the sphere in various mean curvature regimes. The outcomes we obtain are different from those previously known on compact manifolds without CKVF. The first distinction arises from existence/nonexistence of solutions in the near-constant mean curvature regime as we will show below. For manifolds having no CKVF, previous works such as [25, 31, 11] have shown that the reduced system (1.7) admits a unique solution $\varphi$ whenever $\rho\not\equiv 0$ and the ratio $\frac{d\tau}{\tau}$ is sufficiently small. Solutions in this case are the so-called near-CMC solutions, and their existence may be obtained through several approaches. However, in the theorem below, we will see that this conclusion no longer holds when the manifold has CKVF, with the following setting on the round sphere as a counterexample.
@@
which coincides with the conformal equations (1.2) of null TT-tensor $\sigma$ in the vacuum case. The question of existence of solutions to these equations was first posed by D. Maxwell in [33] and it was shown by the second author in [36, 37] to be central to the understanding of the full vacuum case of (1.2). In the absence of CKVF, the second author also proved in [36] that if $\tau>0$ satisfies
@@
  • When $\mathcal{B}_{\tau,\psi}\leq 0$, that is the defocusing case, the system is known to be stable in the CMC and near-CMC settings thanks to well-known arguments in [3, 31, 11]. However, when $\tau$ is far–from–CMC, the stability question still remains open.
@@
For $\mathcal{B}_{\tau,\psi}\leq 0$, we focus on a simplified setting in which $(V,\psi,\sigma)$ vanish identically, that is the equations (1.7). Using the condition $S_{V,\psi,\rho,\varphi}\geq 0$, we show that, within the space of radial functions, the equations are not only stable, but that the set of solutions is also uniformly bounded, independently of the choice of $\tau$. This suggests that the stability property of the conformal equations may persist more generally in this case. When $\mathcal{B}_{\tau,\psi}$ changes sign, the situation is markedly different. In contrast to the stability result obtained by O. Druet and E. Hebey [16] for the CMC case in dimensions $n\leq 5$ as mentioned above, thanks to Identity (1.6), we will construct a sequence of blowing-up solutions to (1.2), thereby showing that the stability property of the conformal equations no longer holds in general in the non-CMC setting. This gives a negative answer to the stability question in the last case in all dimensions. More precisely, our associated results are that
Main Theorem 5 (Non-positive $\mathcal{B}_{\tau,\psi}$ and stability).
@@ 1.2. Hyperbolic manifold
Then the conformal equations (1.7) associated with $(\rho,\tau)$ admit at least one radial solution. Moreover, the set of solutions is compact.
@@ 1.3. Euclidean manifold
In this article, by restricting to radial settings, we do not encounter these limitations, as we can apply Main Theorem 1 to solve (1.7) on the Euclidean space. Our main result is that, for any radial data set, the equations are always solvable, and moreover, the set of solutions is uniformly bounded independently of $\tau$.
@@
There exists a vacuum constraint solution $(\mathbb{R}^{n},g,k)$ satisfying $g-\delta_{\text{Euc}}=O\big(\frac{1}{r^{n-2}}\big)$ and $k=O\big(\frac{1}{r^{n/2}}\big)$ at infinity, and such that the ADM mass of $(\mathbb{R}^{n},g)$ is negative.
@@
In Section 3, provided that the manifold is harmonic, building on the result in [34], we will reduce the conformal equations (1.2) of radial settings to a single nonlinear equation involving functions of the distance, which is much easier to solve.
In Sections 4, 5 and 6, a complete description of this single equation will be obtained in the special cases where the manifold is the sphere, the hyperbolic space or the Euclidean one. Each of these sections begins with the proof of our Main Theorem 1 in the setting considered, and then derives applications by studying the properties of solutions.
@@ 2.1. Preliminaries
There is a rich literature on the existence and classification of harmonic manifolds. The most basic examples are Euclidean spaces and the Rank One Symmetric Spaces (ROSS). In 1944, A. Lichnerowicz conjectured — proving it in dimension 4 — that these are the only harmonic manifolds. The conjecture was later confirmed by Z.I. Szabó for compact, simply connected manifolds of any dimension (see [43]). However, in a significant development, E. Damek and F. Ricci showed in [12] that the conjecture fails in the non-compact case by constructing harmonic homogeneous manifolds that are not ROSS. To date, the only known harmonic manifolds are the Euclidean spaces, ROSS, and Damek–Ricci spaces. For an introduction to the basic structure of ROSS, we refer the reader to [5, Chapter 3] or [29]. General properties of harmonic manifolds can be found in [5, chapter 6], [43, sections 1 and 2] and [28, section 2].
@@
Point (a) is a standard result of Riemannian geometry. For point (b), see [28, Corollary 2.8 and Section 6]. Point (c) follows from Szabó’s theorem and the standard properties of compact ROSS in [43]. ∎
@@ 2.2. Primary functions for the conformal equations
[…]
@@
Assume that $(M,g)$ satisfies the one of two conditions (2.8). For any $f_{v}$ given as in (2.9), we define $F_{v}$ as follows.
@@ 3.2. Reduction of the conformal equations
$V\in C^{0}(\mathbb{R}),\quad\psi(r)\in C^{1}(M),\quad\tau(r)\in\ C^{1}(M)\quad\text{and}\quad\rho(r)\in C^{0}(M).$
@@ 4.1. Lichnerowicz-type equation on the sphere
By equation (2.12) defining the function $F_{v}$, and from (4.1) we have
@@ 4.2. Non-existence of solutions
In the following result, we will give a different perspective on the conformal equations on the sphere. In fact, in the presence of CKVF makes the above results are no longer valid. Otherwise, it would lead to an unexpected situation where the set of solutions to the vacuum conformal equations with null $\sigma$ is infinite. To see this, we first remark that the set of all radial mean curvatures $\tau$ on the sphere well covers the non-generic inequality considered in [36, 37], that is, there exists a class of radial mean curvatures satisfying (4.13). A simple example of this is $\tau_{0}=\cos(r)+2$. By straightforward calculations, since
@@ 4.3. Existence of solutions
    $\sup_{a\in[-c,c]}\bigg(\frac{\max|(\tau_{0}-a\tau_{1})^{\prime}|}{\min|(\tau_{0}-a\tau_{1})|}\bigg)<\frac{2}{\pi}$
@@
$\sup_{a\in[-c,c]}\Big(\frac{\max|(\tau_{0}-a\tau_{1})^{\prime}|}{\min|(\tau_{0}-a\tau_{1})|}\Big)<\frac{2}{\pi},$
@@
has no nontrivial $W$ for all $\lambda\in(0,1]$, then the vacuum conformal equations (1.2) admit at least one solution $(\varphi,W)$. Moreover, in this case, the set of solutions to (1.2) is compact. The results in [36, 37, 15] also show that the limit equation (4.36) appears to play a central role in the existence of solutions to the conformal equations, at least for $\tau$ not changing sign. Technically, as we may see in the proof of the main result in [11] or [35], a nontrivial solution $W$ to the limit equation (4.36) arises from a blow-up phenomenon of a sequence of solutions $\{\varphi_{i}\}$, for which, after an appropriate rescaling, one obtains
$\bigg(\frac{\varphi_{i}}{\|\varphi_{i}\|_{L^{\infty}}}\bigg)^{N}\longrightarrow\sqrt{\tfrac{n}{n-1}}\frac{|LW|}{\tau}.$ (4.37)
Therefore, in our setting of a radial data set and considering radial solutions, we have that the functions $\varphi_{i}$ and $\tau$ in (4.37) are radial. Therefore, the convergence tells us that $|LW|$ is also radial, and hence by Proposition 3.3, $W$ must be of the form $W=u(r)dr$, for some radial function $u(r)$. However, in this case, the following proposition shows that such a non-trivial solution $W$ does not exist.
@@
Although the limit equation criterion cannot be applied for the conformal equations on the sphere due to the presence of CKVF, this proposition provides insight that the equations (4.12) seems to be solvable in some sense. In the next two results, we will give existence of solutions to (4.12) in the far–from–CMC regime by generalizing the results of Theorem 4.7, in which the conditions requiring mean curvature to be near constant and time-derivative $\rho$ to be small may be dropped; instead, all we need to assume is that $(\tau_{0}+\alpha\tau_{1})$ does not change sign or that $\rho$ is $C^{1}$-regular. More precisely, our results are as follows.
@@
an analysis similar to that in [11, Lemma 2.6] or [35, Theorem 3.3] shows that (up to a subsequence)